### v1

@@ 4.5 Reasoning about enforced guidelines
In many cases, protocols or published guidelines for allocating treatment have been established. Physicians may not follow the guidelines in all cases, and so this may differ from the SOC, even if the guidelines were available during the period the data were generated. Determining if Physicians’ implementation of the guidelines of the SOC during a period when the guidelines were being used, and enforcement of guidelines may be of interest. This provides information about how closely the protocol has been followed and, more importantly, if unmeasured covariates are used to improve the outcomes over the guidelines. This gap should always be considered before replacing current guidelines with a new treatment regime. This, of course, is only possible if guidelines were available during the period while the data were generated, which is not the case in our setting. It is, however, also of interest to see if the guidelines, or possibly just time, improved outcomes over the observable SOC. Thus, one should compare both the estimated optimal regime and SOC to enforcing the current clinical guidelines.
However, as few, if any, datasets will indicate an indicator that the physician did or did not follow the guidelines, this requires translation of the often vague guidelines to the functions of the observed data. This translation will be contextual and as mentioned above can only be based on observed variables.
The EULAR guidelines are the guidelines for switching to second line treatment within RA [9]. They consist of five overarching principles and ten specific recommendations. To translate these guidelines into a treatment regime $f_{cgl}$, we have chosen to categorize medications into groups that align with those specified in the guidelines. Therefore, we focus on interventions based on treatment groups rather than individual drugs. The guidelines outline five primary treatment groups: TNF inhibitors, interleukin inhibitors, JAK inhibitors, conventional synthetic (sc) DMARDs, and glucocorticoids (GC). For reasons unknown to us, in the Danish register data no individual has been treated with JAK inhibitors (violating the positivity assumption) and hence this treatment group is not allocated in the $f_{cgl}$ even though present in the guidelines. Some recommendations propose to make a decision based on the present of poor prognostic factors, which is defined through a list of indicators. We have then used these indicators and cross referenced them with the available observations in the Danish register data. With this we can define $f_{cgl}$:
If pre-treatment-switch CRP $<$ 10 mg/L (indicative of poor prognosis factors): Treat with csDMARD +/- GC (which we label csDMARD). But if pre-treatment-switch CRP $\geq$ 10 mg/L: Treat with biologics +/- csDMARD (which we label biologics).
@@ 4.6 Estimated optimal regime
The estimated optimal DTR, denoted by $f_{opt}$, is estimated using a machine learning approach that incorporates random Fourier features (RFF). This method estimates the best treatment decisions based on measured covariates from a random hold-out training sample, ensuring generalizability. The model has been trained on all available covariates in the Danish register data sample (Table 2).
@@ 4.7 Results
Figure 1: Clinical utilities with 95% confidence intervals for the three DTRs.
Figure 1 shows the clinical utilities with their 95% confidence intervals, estimated using both the binary g-computation and the binary IPW estimator. As this is a real world example we cannot guarantee to have observed all confounders. Similarly, the propensity score model and the outcome regression model might be wrongly specified. This is the most likely reason for the different confidence intervals between the bootstrapped and the estimated confidence intervals (supplementary material section 2.2) as well as to different estimands between the IPW and g-computation estimators.
@@
This conclusion is not surprising as our estimated optimal regime was derived simply, for demonstrative purposes, and is unlikely to be superior to guidelines put into place by a group of experts. However, it may be somewhat surprising that even a simple estimated optimal regime would do no better than the SOC, which is a important illustration of why the comparison to the SOC is essential for determining if an estimated optimal regime should be considered for use in practice.
As noted above in Subsection 4.4, the guidelines are not fully enforceable from the available data, as there will always be room for physician’s input within the treatment decision process. This also illustrates why one might always wish to compare the SOC to the part of the guidelines that are enforceable via the available data, as this speaks to potentially unmeasurable physician’s input improving on the guidelines.
As the guidelines were not published during data accrual, our comparison between the enforceable guidelines and the SOC does not provide such insights. It would, therefore, be of scientific interest to redo this analysis after the guidelines have been in use for a number of years to determine how close enforcing the guidelines is to the observed SOC.
@@ 5 Discussion
It is crucial to medical decision-making to optimize treatment strategies based on current information and potential interventions. Although this is the target of the DTR methods, less attention has been paid to comparisons of optimal treatment decisions to SOC or accounting for physician’s ability to make better decisions than the data alone would allow. Sachs et al.[3] suggest the estimand of clinical utility, which compares an estimated optimal decision rule to the SOC. By integrating the standard of care into the evaluation of treatment strategies, it can be easily determined whether it might be worthwhile to attempt implementing novel data-driven DTRs or if the SOC already achieves similar performance. We see comparing any ‘optimal’ decision rule to the SOC and the current guidelines as an alternative method to including the physician’s decision in the optimization, as suggested by Stensrude et al.[19].
@@
We have outlined several practical details of the sample selection and estimation procedures not highlighted in Sachs et al.[3]. In particular, the enforcement and comparison to the current guidelines were not considered in [3], which we show in the real-data example to be valuable information. Additionally, we outline the consideration for sample selection for determining the subjects that should be considered withing the eligible population for the standard of care. Finally, in the real data example we demonstrate how one can select a time zero for the emulated trial comparing the optimal decision rule to the SOC or enforced guidelines.
We have restricted ourself exclusively to use a discrete random variable $T$ for the assigned treatment. Future research could extend this work by investigating the use of continuous $T$ as well as an extension for intervening at multiple time points.
Our hope is that this work both understandably provides the methods and demonstrates their use to underscore the importance of integrating the standard of care into treatment optimization processes to enhance clinical decision-making and ultimately improve patient outcomes. Future research exploring alternative methodologies and other real-world applications of the proposed estimators may further advance our understanding and application of clinical utility in medical practice.

### v2

@@ 4.4 Reasoning about standard of care
If the SOC yields better outcomes than a regime optimized over measured covariates, it implies one of two possibilities: either the estimated optimal is not actually optimal (e.g., due to finite sample bias or model misspecification) or physicians base their decisions on unmeasured prognostic factors. In the latter case, the SOC uses information unavailable to the researcher, which can make it superior to any rule derived solely from the observed data. However, also in this setting the clinical utility is not identified based on $Z$ alone.
@@
It is important to note that the standard of care may be less stable over time than other covariates. As highlighted by [18], the distribution of treatment decisions made by physicians especially when influenced by evolving guidelines can shift in response to new evidence. This instability makes it more difficult to generalizing clinical utility estimates across time or settings. In contrast, biological covariates such as biomarkers tend to be more stable. Therefore, when evaluating the clinical utility of a decision rule relative to the SOC, it is important to consider the possibility of temporal drift in physician behavior, which may affect the observed outcomes.
@@ 4.5 Reasoning about enforced guidelines
In many cases, protocols or published guidelines for allocating treatment have been established. Physicians may not follow the guidelines in all cases, and so this may differ from the SOC, even if the guidelines were available during the period the data were generated. Determining if Physicians’ implementation of the guidelines of the SOC during a period when the guidelines were being used, and enforcement of guidelines may be of interest. This provides information about how closely the protocol has been followed and, more importantly, if the standard of care improves the outcomes over the guidelines. This gap should always be considered before replacing current guidelines with a new treatment regime. This, of course, is only possible if guidelines were available during the period while the data were generated, which is not the case in our setting. It is, however, also of interest to see if the guidelines, or possibly just time, improved outcomes over the observable SOC. Thus, one should compare both the estimated optimal regime and SOC to enforcing the current clinical guidelines.
However, as few, if any, datasets will include an indicator that the physician did or did not follow the guidelines, this requires translation of the often vague guidelines to the functions of the observed data. This translation will be contextual and as mentioned above can only be based on observed variables.
The EULAR guidelines are the guidelines for switching to second line treatment within RA [9]. They consist of five overarching principles and ten specific recommendations. To translate these guidelines into a treatment regime $f_{cgl}$, we have chosen to categorize medications into groups that align with those specified in the guidelines. Therefore, we focus on interventions based on treatment groups rather than individual drugs. The guidelines outline five primary treatment groups: TNF inhibitors, interleukin inhibitors, JAK inhibitors, conventional synthetic (sc) DMARDs, and glucocorticoids (GC). For reasons unknown to us, in the Danish register data no individual has been treated with JAK inhibitors (violating the positivity assumption) and hence this treatment group is not allocated in the $f_{cgl}$ even though present in the guidelines. This also implies that our results cannot be directly generalized to settings where JAK inhibitors are available. Some recommendations propose to make a decision based on the present of poor prognostic factors, which is defined through a list of indicators. We have then used these indicators and cross referenced them with the available observations in the Danish register data. With this we can define $f_{cgl}$:
If pre-treatment-switch CRP << 10 mg/L (indicative of poor prognosis factors): Treat with csDMARD +/- GC (which we label csDMARD). But if pre-treatment-switch CRP $\geq$ 10 mg/L: Treat with biologics +/- csDMARD (which we label biologics).
@@ 4.6 Estimated optimal regime
The estimated optimal DTR, denoted by $f_{opt}$, is estimated using a machine learning approach that incorporates random Fourier features (RFF). This method estimates the best treatment decisions based on measured covariates from a random hold-out training sample, ensuring generalizability. The estimated optimal regime should be interpreted as optimal within the space of measured covariates (L-optimal). If unmeasured factors modify treatment effects, the true optimal regime may differ, and the estimand may not be identified. The model has been trained on all available covariates in the Danish register data sample (Table 2).
@@ 4.7 Results
Figure 1: Estimated clinical utilities with 95% confidence intervals for the three DTRs.
Figure 1 shows the clinical utilities with their 95% confidence intervals, estimated using both the binary g-computation and the binary IPW estimator. We used standard logistic regression implementations for both the outcome and propensity score models. As this is a real world example we cannot guarantee to have observed all confounders. Similarly, the propensity score model and the outcome regression model might be wrongly specified. This is the most likely reason for the different confidence intervals between the bootstrapped and the estimated confidence intervals (Supplementary Material Section 2.2) as well as to different estimands between the IPW and g-computation estimators. The ‘estimated’ confidence intervals are derived analytically using M-estimation theory (sandwich estimator), which relies on asymptotic normality (see Supplementary Material Section 2.1). In contrast, the bootstrapped intervals are obtained by resampling the dataset.
@@
This conclusion is not surprising as our estimated optimal regime was derived simply, for demonstrative purposes, and is unlikely to be superior to guidelines put into place by a group of experts. However, it may be somewhat surprising that even a simple estimated optimal regime would do no better than the SOC, which is a important illustration of why the comparison to the SOC is essential for determining if an estimated optimal regime should be considered for use in practice. Interestingly, the guidelines, which are also a function of measured covariates, appear superior to the estimated optimal regime. The estimated optimal regime is derived by optimizing a prediction model (i.e., minimizing a prediction loss), which is distinct from maximizing the clinical utility. Due to this distinction, together with likely finite sample bias and potential model misspecification, the estimated regime is not truly optimal for the target estimand.
As noted above in Subsection 4.4, the guidelines are not fully enforceable from the available data, as there will always be room for physician’s input within the treatment decision process. This also illustrates why one might always wish to compare the SOC to the part of the guidelines that are enforceable via the available data, as this speaks to physicians improving on the guidelines. As the guidelines were not published during data accrual, our comparison between the enforceable guidelines and the SOC does not provide such insights. It would, therefore, be of scientific interest to redo this analysis after the guidelines have been in use for a number of years to determine how close enforcing the guidelines is to the observed SOC.
@@ 5 Discussion
It is crucial to medical decision-making to optimize treatment strategies based on current information and potential interventions. Although this is the target of the DTR methods, less attention has been paid to comparisons of estimated optimal DTR to the SOC. Sachs et al.[3] suggest the estimand of clinical utility, which compares an estimated optimal decision rule to the SOC. By integrating the standard of care into the evaluation of treatment strategies, it can be easily determined whether it might be worthwhile to attempt implementing novel data-driven DTRs or if the SOC already achieves similar performance.
@@
We have outlined several practical details of the sample selection and estimation procedures not highlighted in Sachs et al.[3]. In particular, the enforcement and comparison to the current guidelines were not considered in [3], which we show in the real-data example to be valuable information. Additionally, we outline the consideration for sample selection for determining the subjects that should be considered withing the eligible population for the standard of care. Finally, in the real data example we demonstrate how one can select a time zero for the emulated trial comparing the estimated optimal decision rule to the SOC or enforced guidelines.
As we have seen with the data example, if the SOC differs across populations or evolves over time, the clinical utility may also differ as it is relative to the SOC. Conditions that allow generalization to different target distributions may be possible to derive by adapting methods from the transportability literature [21]. For example, if we assume the target population has the same distributions of treatment options and covariates then we can likely apply our results to this population as well. Instead of attempting to transport the clinical utility, updating the clinical utility of the estimated optimal DTR over time or calculating it in a new setting where the SOC may differ is advisable. Additionally, as with all prediction or machine learning models, data drift has to be considered from the training population in comparison to the target population [22], as this can cause performance deterioration.
This suggests that not only should the clinical utility of an optimal DTR be reevaluated with the changing SOC, but that the DTR should potentially be rederived in new settings and as patients’ characteristics shift within the same population. Since the clinical utility is dependent on the specific SOC of the analyzed population, we recommend explicitly characterizing the observed treatment patterns (such as treatment distributions within subgroups) to aid any future meta-analysis or comparisons.
Our analysis focused on a single treatment decision for clarity and feasibility. Many clinical contexts, however, involve multiple sequential decisions, and extending our framework to such settings would be of great interest. Multi stage regimes introduce additional complexity, including the need to model time varying confounding, ensure sequential exchangeability, and maintain positivity across treatment histories. These challenges are well documented in the literature on dynamic treatment regimes [7]. Furthermore, comparing sustained regimes to the standard of care is complicated by the fact that observed post-baseline treatments may not correspond to the natural course under earlier interventions [23]. Future work should explore these extensions to better capture the longitudinal nature of treatment decisions.
We see comparing any estimated optimal decision rule to the SOC and the current guidelines as an alternative method to including the physician’s decision in the optimization, as suggested by Stensrud et al.[18], or other related works, such as Levis et al.[24], which propose a measure of the potential benefit of a targeted intervention compared to status quo, conditional on covariates, which is closely related to clinical utility. We note that the Levis et al. [24] approach also allows for the incorporation of supply constraints, making it particularly useful in settings where the intervention of interest is expensive or difficult to obtain. Such restrictions could also be included in the derivation of any DTR, prior to evaluation against SOC using clinical utility.
Our hope is that this work both understandably provides the methods and demonstrates their use to underscore the importance of integrating the standard of care into treatment optimization processes to enhance clinical decision-making and ultimately improve patient outcomes. Future research exploring alternative methodologies and other real world applications of the proposed estimators may further advance our understanding and application of clinical utility in medical practice.